### v1

@@ 0.B.1 Fine-grained SugarCrepe performance
The SugarCrepe [33] benchmark comprises of seven compositional manipulations of captions for a subset of MS COCO [41]. The benchmark evaluates the zero-shot accuracy of vision-language models for picking the ground-truth captions over the generated hard negatives. We provide the fine-grained results of our models for each of the seven SugarCrepe manipulations in Tab. 0.B.1. Individual categories often demonstrate trade-offs amongst each other, but the averaged performance reveals consistent trends of Sparo outperforming the baselines. Additionally, we show that the compositionality exposed in learned Sparo encodings across different categories can be enhanced by performing post-hoc concept selection, demonstrated in Sec. 5.1.
0.B.2 Comparison with other bottlenecks
Table 0.B.2: Comparing zero-shot recognition, robustness, and SugarCrepe compositionality for CLIP models with a linear (affine) information bottleneck and a discrete bottleneck using Finite Discrete Tokens (FDT) [13] trained on Conceptual Captions.
Train Model ImageNet- Object Net Sugar Crepe
V1 V2 Sketch A R
CC3M CLIP32 (C) 0.109 0.091 0.053 0.025 0.154 0.064 0.578
$\mathcal{C}$+GAP 0.124 0.105 0.064 0.025 0.178 0.078 0.582
$\mathcal{C}$+Sparo 0.132 0.113 0.072 0.027 0.190 0.083 0.596
CLIP16 (C) 0.141 0.122 0.068 0.033 0.177 0.080 0.611
$\mathcal{C}$+GAP 0.156 0.134 0.069 0.033 0.187 0.081 0.616
$\mathcal{C}$+Sparo 0.170 0.140 0.088 0.035 0.221 0.098 0.625
CC12M CLIP32 (C) 0.304 0.257 0.201 0.060 0.405 0.160 0.691
$\mathcal{C}$+GAP 0.320 0.272 0.221 0.063 0.426 0.178 0.692
$\mathcal{C}$+Sparo 0.344 0.293 0.251 0.069 0.468 0.200 0.693
CLIP16 (C) 0.361 0.311 0.249 0.091 0.467 0.218 0.697
$\mathcal{C}$+GAP 0.382 0.330 0.262 0.101 0.501 0.241 0.695
$\mathcal{C}$+Sparo 0.406 0.350 0.298 0.113 0.559 0.268 0.723
CC15M CLIP32 (C) 0.329 0.279 0.232 0.070 0.435 0.178 0.687
$\mathcal{C}$+GAP 0.345 0.294 0.242 0.077 0.460 0.192 0.678
$\mathcal{C}$+Sparo 0.370 0.313 0.271 0.083 0.510 0.213 0.710
CLIP16 (C) 0.384 0.337 0.268 0.105 0.503 0.238 0.699
$\mathcal{C}$+GAP 0.399 0.343 0.287 0.114 0.531 0.252 0.701
$\mathcal{C}$+Sparo 0.437 0.378 0.317 0.145 0.579 0.279 0.730
Setup.
Linear information bottleneck details.
Discrete bottleneck details.
Results.
We present our results in Tab. 0.B.2. First, we see that a simple linear bottleneck performs worse than standard CLIP in a majority of settings, indicating that the mechanism behind the representational bottleneck is important. As expected, the discrete bottleneck widely outperforms or stays comparable to standard CLIP. However, note that global average pooling (GAP) continues to shine as a strong baseline, outperforming the discrete bottleneck in most settings. Finally, we verify that Sparo outperforms or remains comparable to the best baselines across all settings.
0.B.3 Additional visualizations
@@
Figure 0.B.1: Additional visualizations of attended image and text positions for three Sparo slots (one per row). We surmise that the Sparo concepts from top to bottom represent animals, transportation, and seats. Notice that top-right and center-right examples have similar attention masks over ‘horse,’ but consider different aspects of the concept – one as an animal, another as a mode of transportation.
We present additional examples of learned Sparo concepts in Fig. 0.B.1 following the filtering described in Sec. 5.5. Each row illustrates the attended positions of the single-head attention operation for the same Sparo slot across examples from MS COCO [41]. We reuse the same last input for the first two visualized slots to highlight cases where visually similar attention maps can correspond to distinct concepts. Looking at attended positions allows us to interpret where each slot takes information from, but not how that information is used.

### v2

@@ 0.B.4 Fine-grained SugarCrepe performance
The SugarCrepe [34] benchmark comprises of seven compositional manipulations of captions for a subset of MS COCO [42]. The benchmark evaluates the zero-shot accuracy of vision-language models for picking the ground-truth captions over the generated hard negatives. We provide the fine-grained results of our models for each of the seven SugarCrepe manipulations in Tab. 0.B.3. Individual categories often demonstrate trade-offs amongst each other, but the averaged performance reveals consistent trends of Sparo outperforming the baselines. Additionally, we show that the compositionality exposed in learned Sparo encodings across different categories can be enhanced by performing post-hoc concept selection, demonstrated in Sec. 5.1.
0.B.5 Winoground performance
Table 0.B.4: Winoground zero-shot accuracies for CLIP models trained on Conceptual Captions and LAION-400M.
Train Model Text Image Group
CC3M CLIP16 (C) 0.205 0.075 0.052
$\mathcal{C}$+GAP 0.190 0.097 0.065
$\mathcal{C}$+Sparo 0.215 0.123 0.072
CC12M CLIP16 (C) 0.248 0.087 0.052
$\mathcal{C}$+GAP 0.245 0.095 0.052
$\mathcal{C}$+Sparo 0.245 0.102 0.058
CC15M CLIP16 (C) 0.243 0.085 0.052
$\mathcal{C}$+GAP 0.245 0.078 0.052
$\mathcal{C}$+Sparo 0.273 0.102 0.075
L400M CLIP32 (C) 0.287 0.110 0.083
$\mathcal{C}$+GAP 0.265 0.132 0.087
$\mathcal{C}$+Sparo 0.278 0.140 0.097
We evaluate CLIP+Sparo against CLIP and CLIP+GAP on Winoground [65], and report the metrics suggested by the benchmark. Given image-caption pairs $(I_{0},C_{0})$ and $(I_{1},C_{1})$, the ‘Text’ score averages whether the model can select the correct captions $C_{0}$ and $C_{1}$ given the images $I_{0}$ and $I_{1}$ respectively. The ‘Image’ score tallies whether the model can select the correct images $I_{0}$ and $I_{1}$ for the captions $C_{0}$ and $C_{1}$ respectively. The ‘Group’ score combines these tests, and evaluates whether the model can select both the correct image given the caption, and the correct caption given the image, for combinations of image-text pairs. We report our results in Tab. 0.B.4, showing superior performance of Sparo in all of the ‘Image’ and ‘Group’ evaluations. For ‘Text’, we find that Sparo performs as well as or better than GAP, but CLIP without GAP outperforms both in two out of our four training setups.
0.B.6 Absolute VTAB results
Table 0.B.5: Absolute VTAB benchmark results corresponding to the relative differences illustrated in Fig. 2.
VTAB dataset CC3M ($\mathcal{C}^{16}$) CC12M ($\mathcal{C}^{16}$) CC15M ($\mathcal{C}^{16}$) L400M ($\mathcal{C}^{32}$)
GAP Sparo GAP Sparo GAP Sparo GAP Sparo
cifar100 0.163 0.168 0.330 0.390 0.397 0.449 0.707 0.742
resisc45 0.169 0.184 0.326 0.359 0.357 0.360 0.527 0.563
eurosat 0.193 0.226 0.276 0.335 0.246 0.292 0.480 0.513
dtd 0.104 0.121 0.162 0.234 0.181 0.195 0.502 0.506
svhn 0.074 0.097 0.087 0.098 0.084 0.125 0.236 0.458
caltech101 0.539 0.562 0.731 0.744 0.752 0.748 0.828 0.829
smallnorb_elevation 0.114 0.119 0.100 0.108 0.138 0.116 0.102 0.108
smallnorb_azimuth 0.055 0.053 0.052 0.053 0.058 0.057 0.049 0.053
clevr_closest_obj 0.224 0.194 0.151 0.197 0.109 0.095 0.204 0.082
clevr_count 0.105 0.132 0.227 0.232 0.136 0.151 0.189 0.164
pcam 0.487 0.495 0.477 0.535 0.502 0.572 0.549 0.495
sun397 0.261 0.263 0.464 0.489 0.505 0.519 0.671 0.663
dmlab 0.148 0.172 0.172 0.135 0.134 0.171 0.202 0.141
kitti_closest_dist 0.118 0.322 0.421 0.392 0.249 0.429 0.245 0.125
diabetic_ret 0.066 0.147 0.165 0.118 0.166 0.174 0.047 0.023
dsprites_xpos 0.033 0.028 0.030 0.029 0.035 0.044 0.032 0.030
dsprites_orientation 0.027 0.020 0.029 0.026 0.019 0.015 0.017 0.025
pets 0.099 0.104 0.587 0.575 0.595 0.579 0.860 0.869
flowers 0.119 0.131 0.298 0.287 0.324 0.311 0.659 0.673
In Tab. 0.B.5, we present the absolute numbers behind the relative improvements of Sparo illustrated in Fig. 2 for the VTAB benchmark [78].
0.B.7 Additional visualizations
@@
Figure 0.B.2: Additional visualizations of attended image and text positions for three Sparo slots (one per row). We surmise that the Sparo concepts from top to bottom represent animals, transportation, and seats. Notice that top-right and center-right examples have similar attention masks over ‘horse,’ but consider different aspects of the concept – one as an animal, another as a mode of transportation.
We present additional examples of learned Sparo concepts in Fig. 0.B.2 following the filtering described in Sec. 5.4. Each row illustrates the attended positions of the single-head attention operation for the same Sparo slot across examples from MS COCO [42]. We reuse the same last input for the first two visualized slots to highlight cases where visually similar attention maps can correspond to distinct concepts. Looking at attended positions allows us to interpret where each slot takes information from, but not how that information is used.